# Néel vector controlled charge and spin transport in altermagnetic junctions

### v2

@@ Abstract
We theoretically investigate electron transport in junctions between the two altermagnets (AMs) in strong and weak altermagnetic phases. The charge and spin conductivities are analyzed as functions of angle $\theta$ between the Néel vectors of the two AMs. In the strong AM regime, the charge conductivity vanishes as $\theta\to\pi$, while in the weak AM regime it remains finite. Introducing a normal metal between two AMs leads to Fabry–Pérot-type oscillations in charge conductivity which can be controlled by an applied gate voltage. In the strong regime, transport is dominated by up-spin electrons, whereas both spin channels contribute in the weak regime. These results highlight the potential of AM-based heterostructures for spintronic applications, such as spin filters, and quantum interference–based spintronic devices, where tunable spin-dependent transport and interference effects can be utilized in electronic devices.
@@ I Introduction
Altermagnets (AMs), materials having $d$-wave magnetic order have generated tremendous interest among condensed matter physicists in the past couple of years [smejkal22a, smejkal22b, smejkal22c, fern24, zhou24, yan24, fu2025am, yu25]. Characterized by traits of both ferromagnets and antiferromagnets, their net spin polarization is zero. They are known to carry spin current under a voltage bias [das2023]. Junctions of AMs with normal metals, ferromagnets and superconductors have been studied by many groups [das2023, papaj, sun23, das2024car, fu2025]. The spin that commutes with the Hamiltonian of AM defines the Néel vector for the AM. Several candidate materials such as MnTe, Mn5Si3, KV2Se2O exist for AMs [Helena21, Jiang2025, Amin2024].
Magnetic tunnel junctions (MTJs) are junctions between two ferromagnetic metals separated by a thin insulator, wherein electrons are able to pass through the thin insulating barrier. The relative orientation of the magnetic moments in these layers dictates the possibility of electron tunneling. The parallel alignment of magnetizations in the ferromagnetic layers exhibits a low electrical resistance in the junction, whereas an antiparallel alignment results in a high resistance. This variation in resistance between the two magnetic configurations gives rise to the tunneling magnetoresistance (TMR) effect, which forms the fundamental basis for the operation of MTJ-based spintronic devices [Julliere, miyazaki1995, Yuasa2007, Moodera95]. Magnetic tunnel junctions in altermagnetic RuO2, and single ferromagnetic electrode have been shown to result in a high tunneling magnetoresistance [Jiang23, Noh25, Samanta24]. Magnetic tunnel junctions between the altermagnets has just recently been studied where the authors claim to achieve tunneling magnetoresistance over 1000% by just rotating the AM and tuning the altermagnetic strength and Fermi energy [Sun25, Ezawa25].
The orientation of the Néel vector can be tuned using spin–orbit torques or ultrafast optical excitations [Zhang22, Ono2021]. It influences how spin-polarized currents propagate through the materials. A recent study on AM/$p$-wave magnet junction by rotating the Néel vector of the latter relative to that of the former shows similar response [das25]. In MnTe, it is found that domains which have different directions for the Néel vectors are formed very much like that in ferromagnets [Amin2024].
@@
In Sec. II, we present the general outline of our calculation. Sections III and IV discuss AM–AM and AM–NM–AM junctions in the weak AM regime. Sections V and VI cover the same junction types in the strong AM regime. Each section from III to VI is divided into two parts. The first part contains the detailed calculations, and the second part presents the corresponding results with analysis. Finally, we summarise our findings and provide concluding remarks in Sec. VII.
@@ II Outline of calculation
where $a$ is the lattice spacing, $\sigma_{j}$’s are Pauli spin matrices. By Taylor expanding such a Hamiltonian in momentum space near the band bottom, its continuum form can be obtained. Depending on the relative strength of the altermagnetic term compared to the normal metal term in the Hamiltonian, the phase of AM can be classified into strong and weak. The strong phase corresponds to $t_{J}>t\geq 0$ whereas the weak phase corresponds to $0\leq t_{J}<t$.
@@ III.1 Details of the calculation
Here, $c$ and $q_{0}$ characterise the junction. In certain limits, $c$ can be thought of physically as the ratio between the hopping strength at the bond that forms the junction to $t$, and $q_{0}$ corresponds to the strength of delta-function impurity at the junction [das2023].
@@
$G^{s\pm}$ $=$ $\frac{e}{8\pi^{2}\sqrt{t^{2}-t_{J}^{2}}}\int_{-\pi/2}^{\pi/2}d\alpha[J^{s\pm}_{\uparrow}(\alpha)+J^{s\pm}_{\downarrow}(\alpha)]$ (11)
III.2 Results
In Fig. 1(b,c), we plot the conductivities versus the angle between the Néel vectors of the two AMs. In Fig. 1(b), the charge conductivity is shown wavevector the angle $\theta$ for different ratios of $t_{J}/t$, with parameters $q_{0}=1/a$, $c=1$, and $E=t$. For larger values of the altermagnetic strength $t_{J}$, the conductivity exhibits significant variation. We see [Fig. 1(b)] that for $\theta=0$, the charge conductivity is maximum and then decreases monotonically as $\theta$ increases in the range $0,\pi]$. As $\theta$ deviates away from 0, the orientations of the spins corresponding to the same values of $\vec{k}$ on either sides of the junction is different leading to reduced conductivity. Interestingly, the conductivity at $\theta=0$ increases with increasing value of $t_{J}$. This feature can be understood by taking the limit where all the reflection coefficients are zero and using Eq. (11).
@@ IV.1 Details of the calculation
Dispersion for $\uparrow$- and $\downarrow$ spin electrons in the AM is given in Eqn. 3. But the dispersion of the normal metal is given by
@@
The differential charge and spin conductivities in this system are given by Eq.(11).
@@ IV.2 Results
Such oscillations originate from quantum‐interference effects: multiple reflections within the finite NM segment generate constructive or destructive interference determined by the accumulated phase, known as Fabry-Pérot interference(FPI) [soori12, liang2001] . As $L$ increases, the phase acquired by the electron wavefunction varies, leading to alternating constructive and destructive interference. FPI condition given by $\Delta L=\pi/q$, where $\Delta L$ is the interval between successive peaks and $q$ is the wave number of the interfering mode in the NM. This condition is obtained by considering the normal incident electrons which are the dominant contributions to conductivity. $\Delta L$ calculated by the above condition is 2.221a, in agreement with the numerically obtained value of 2.212a obtained in the results in Fig. 2(b).
@@ V.1 Details of the calculation
The differential charge and the spin conductivities are given by Eqn. (11)
V.2 Results
@@
Figure 3(b) shows the variation of charge conductivity with the spin polarization angle $\theta$ for different values of $t$. In the strong phase, conductivity is dominated by a single spin channel, and electron transport occurs only when the transverse momentum matches across the two AM regions. As shown in Fig. 3(a), an electron incident with a given spin is transmitted into the right AM with the same spin. So, there is no down-spin current for the up-spin electron incidence and vice versa. The conductivity is maximum at $\theta=0$ and decreases monotonically with increasing $\theta$, vanishing at $\theta=\pi$. This behavior arises because, for $\theta=0$, the spin orientations in both regions are collinear, enabling large transmission. With increasing $\theta$, the spin overlap is continuously reduced, thus suppressing the conductivity. At $\theta=\pi$, the spins are fully antiparallel, eliminating overlap and hence blocking transmission into the right AM, resulting in zero conductivity.
@@ VI.1 Details of the calculation
The charge- and spin- current densities on two sides of the junction due to this wavefunction are given by Eq. (27), where $J^{s-}_{\downarrow}(\alpha)$ and $J^{s+}_{\downarrow}(\alpha)$ are the spin currents for the left and right AM respectively. The differential charge and spin conductivities in this system are calculated using Eq.(11).
@@ VI.2 Results
VII Summary and conclusion
@@
Recent theoretical proposals have outlined mechanisms to control the Néel vector in altermagnets [han25, zhang25]. These advances indicate that the phenomena we predict may soon be experimentally testable and could play a role in guiding the development of future spintronic device architectures.

### v3

@@ Abstract
Altermagnets (AMs) - magnetic materials that have spin-split bandstructure with zero net spin polarization can be classified as weak or strong depending upon the strength of altermagnetic term in the Hamiltonian. We theoretically investigate electron transport in junctions between the two AMs in strong and weak altermagnetic phases. The charge and spin conductivities are analyzed as functions of angle $\theta$ between the Néel vectors of the two AMs. In the strong AM regime, the charge conductivity vanishes as $\theta\to\pi$, while in the weak AM regime it remains finite. Introducing a normal metal (NM) between two AMs leads to Fabry–Pérot-type oscillations in charge conductivity which can be controlled by an applied gate voltage. In the strong regime, transport in AM-NM-AM junctions is dominated by up-spin electrons, whereas both spin channels contribute in the weak regime. These results highlight the potential of AM-based heterostructures for spintronic applications, such as spin filters, and quantum interference–based spintronic devices, where tunable spin-dependent transport and interference effects can be utilized in electronic devices without a need for externally applied magnetic field.
@@ I Introduction
Altermagnets (AMs), materials having $d$-wave magnetic order have generated tremendous interest among condensed matter physicists in the past couple of years [21, 22, 23, 6, 32, 27, 8, 28]. Characterized by traits of both ferromagnets and antiferromagnets, their net spin polarization is zero. They are known to carry spin current under a voltage bias [3]. Junctions of AMs with normal metals, ferromagnets and superconductors have been studied by many groups [3, 18, 25, 2, 7]. The spin that commutes with the Hamiltonian of AM defines the Néel vector for the AM. Several candidate materials such as MnTe, Mn5Si3, KV2Se2O exist for AMs [19, 10, 1].
Magnetic tunnel junctions (MTJs) are junctions between two ferromagnetic metals separated by a thin insulator, wherein electrons are able to pass through the thin insulating barrier. The relative orientation of the magnetic moments in these layers dictates the possibility of electron tunneling. The parallel alignment of magnetizations in the ferromagnetic layers exhibits a low electrical resistance in the junction, whereas an antiparallel alignment results in a high resistance. This variation in resistance between the two magnetic configurations gives rise to the tunneling magnetoresistance (TMR) effect, which forms the fundamental basis for the operation of MTJ-based spintronic devices [12, 14, 29, 15]. Magnetic tunnel junctions in altermagnetic RuO2, and single ferromagnetic electrode have been shown to result in a high tunnelling magnetoresistance [11, 16, 20]. Magnetic tunnel junctions between the altermagnets has just recently been studied where the authors claim to achieve tunneling magnetoresistance over 1000% by just rotating the AM and tuning the altermagnetic strength and Fermi energy [26, 5].
The orientation of the Néel vector can be tuned using spin–orbit torques or ultrafast optical excitations [30, 17]. It influences how spin-polarized currents propagate through the materials. A recent study on AM/$p$-wave magnet junction by rotating the Néel vector of the latter relative to that of the former shows similar response [4]. In MnTe, it is found that domains which have different directions for the Néel vectors are formed very much like that in ferromagnets [1].
@@
In Sec. II, we present the general outline of our calculation. Sections III and IV discuss AM–AM and AM–NM–AM junctions in the weak AM regime. Sections V and VI cover the same junction types in the strong AM regime. Sections III and V contain three parts. In the first part, we give details of calculations, followed by analytical expressions for reflection amplitudes which will be used in calculating the conductivities, and end with a subsection on results and analysis. Sections IV and VI are divided into two parts. The first part contains the detailed calculations, and the second part presents the corresponding results with analysis. In section VII, we compare our work with some closely related works, followed by a discussion of connection to TMR and the effect of disorder and edge roughness. Finally, we summarise our findings and provide concluding remarks in Sec. VIII.
@@ II Outline of calculation
where $a$ is the lattice spacing, $\sigma_{j}$’s are Pauli spin matrices. Here, $t$ is the hopping strength which characterizes a metal, and $t_{J}$ is spin- and direction- dependent hopping which characterizes altermagnet. By Taylor expanding such a Hamiltonian in momentum space near the band bottom, its continuum form can be obtained. Depending on the relative strength of the altermagnetic term compared to the normal metal term in the Hamiltonian, the phase of AM can be classified into strong and weak. The strong phase corresponds to $t_{J}>t\geq 0$ whereas the weak phase corresponds to $0\leq t_{J}<t$. In the weak regime, the Fermi contours for the two spins intersect whereas in the strong phase, they do not.
@@
Charge conservation ensures that the current is continuous across the junction, providing the basis for formulating appropriate boundary conditions at the interface. These boundary conditions can be mapped to the terms in the Hamiltonian of the underlying lattice Hamiltonian in the vicinity of the junction. Importantly, the resulting boundary conditions differ in the strong and weak altermagnetic regimes. A detailed discussion of the boundary conditions for both cases is presented in the following sections.
@@ III.1 Details of the calculation
Here, $c$ and $q_{0}$ characterise the junction. In certain limits, $c$ can be thought of physically as the ratio between the hopping strength at the bond that forms the junction to $t$, and $q_{0}$ corresponds to the strength of delta-function impurity at the junction [3].
@@
$G^{s\pm}$ $=$ $\frac{e}{8\pi^{2}\sqrt{t^{2}-t_{J}^{2}}}\int_{-\pi/2}^{\pi/2}d\alpha[J^{s\pm}_{\uparrow}(\alpha)+J^{s\pm}_{\downarrow}(\alpha)]$
III.2 Analytical expressions in the limit $\theta=0$
In the limit $\theta=0$, it is straightforward to show that $r_{\uparrow\downarrow}=r_{\downarrow\uparrow}=0$ and
$r_{\uparrow\uparrow}$ $=$ $\frac{tq_{0}+ik_{x\uparrow}(t-t_{J})(1-1/c^{2})}{ik_{x\uparrow}(t-t_{J})(1+1/c^{2})-tq_{0}}$
$r_{\downarrow\downarrow}$ $=$ $\frac{tq_{0}+ik_{x\downarrow}(t+t_{J})(1-1/c^{2})}{ik_{x\downarrow}(t+t_{J})(1+1/c^{2})-tq_{0}}$ (13)
III.3 Results
In Fig. 1(b,c), we plot the conductivities versus the angle between the Néel vectors of the two AMs. In Fig. 1(b), the charge conductivity is shown wavevector the angle $\theta$ for different ratios of $t_{J}/t$, with parameters $q_{0}=1/a$, $c=1$, and $E=t$. For larger values of the altermagnetic strength $t_{J}$, the conductivity exhibits significant variation. We see [Fig. 1(b)] that for $\theta=0$, the charge conductivity is maximum and then decreases monotonically as $\theta$ increases in the range $0,\pi]$. As $\theta$ deviates away from 0, the orientations of the spins corresponding to the same values of $\vec{k}$ on either sides of the junction is different leading to reduced conductivity. Interestingly, the conductivity at $\theta=0$ increases with increasing value of $t_{J}$. This feature can be understood by taking the limit where all the reflection coefficients are zero and using Eq. (LABEL:eq:cond_wk).
@@ IV.1 Details of the calculation
Dispersion for $\uparrow$- and $\downarrow$ spin electrons in the AM is given in Eq. 4. But the dispersion of the normal metal is given by
@@
The differential charge and spin conductivities in this system are given by Eq.(LABEL:eq:cond_wk).
@@ IV.2 Results
Such oscillations originate from quantum‐interference effects: multiple reflections within the finite NM segment generate constructive or destructive interference determined by the accumulated phase, known as Fabry-Pérot interference(FPI) [24, 13] . As $L$ increases, the phase acquired by the electron wavefunction varies, leading to alternating constructive and destructive interference. FPI condition given by $\Delta L=\pi/q$, where $\Delta L$ is the interval between successive peaks and $q$ is the wave number of the interfering mode in the NM. This condition is obtained by considering the normal incident electrons which are the dominant contributions to conductivity. $\Delta L$ calculated by the above condition is 2.221a, in agreement with the numerically obtained value of 2.212a obtained in the results in Fig. 2(b).
@@ V.1 Details of the calculation
The differential charge and the spin conductivities are given by Eq. (LABEL:eq:cond_wk)
V.2 Analytical expressions in the limit $\theta=0$
In the limit $\theta=0$, the reflection amplitudes $r_{\uparrow\downarrow}=r_{\downarrow\uparrow}=0$ and
$r_{\uparrow\uparrow}$ $=$ $\frac{[q_{0}t+i(t_{J}-t)(k_{x\uparrow}-\pi/a)(1-1/c^{2})]}{[-q_{0}t+i(t_{J}-t)(k_{x\uparrow}-\pi/a)(1+1/c^{2})]}$
$r_{\downarrow\downarrow}$ $=$ $\frac{[q_{0}t+ik_{x\downarrow}(t_{J}+t)(1-1/c^{2})]}{[-q_{0}t+ik_{x\downarrow}(t_{J}+t)(1+1/c^{2})]}$ (30)
V.3 Results
@@
Mn5Si3 is a candidate material for AM in the strong regime with $t_{J}=2t=~150meV$ [19]. Figure 3(b) shows the variation of charge conductivity with the spin polarization angle $\theta$ for different values of $t$. In the strong phase, conductivity is dominated by a single spin channel, and electron transport occurs only when the transverse momentum matches across the two AM regions. As shown in Fig. 3(a), an electron incident with a given spin is transmitted into the right AM with the same spin. So, there is no down-spin current for the up-spin electron incidence and vice versa. The conductivity is maximum at $\theta=0$ and decreases monotonically with increasing $\theta$, vanishing at $\theta=\pi$. This behavior arises because, for $\theta=0$, the spin orientations in both regions are collinear, enabling large transmission. With increasing $\theta$, the spin overlap is continuously reduced, thus suppressing the conductivity. At $\theta=\pi$, the spins are fully antiparallel, eliminating overlap and hence blocking transmission into the right AM, resulting in zero conductivity.
@@ VI.1 Details of the calculation
The charge- and spin- current densities on two sides of the junction due to this wavefunction are given by Eq. (29), where $J^{s-}_{\downarrow}(\alpha)$ and $J^{s+}_{\downarrow}(\alpha)$ are the spin currents for the left and right AM respectively. The differential charge and spin conductivities in this system are calculated using Eq.(LABEL:eq:cond_wk).
@@ VI.2 Results
VII Discussion
Transport across AM-NM interfaces has been shown to generate spin currents [3]. Moreover, spin-resolved transport and spin-valve effects have been demonstrated in NM-AM-NM junctions [8]. In contrast, the present work focuses on two different junction geometries: (i) AM-AM and (ii) AM-NM-AM, and systematically explores how charge and spin transport depend on the altermagnetism and the relative orientation of the Néel vectors.
TMR defined as the ratio of the difference between the conductivities in parallel and antiparallel alignment of the two AMs to the smaller of the two (see ref. [26] for this definition) shoots up to infinity for the strong phase since the conductivity in the antiparallel configuration is zero and that in the parallel configuration is finite. We find that TMR in the weak phase turns out to be of the order of $4-12\%$ for AM-NM-AM junctions.
Realistic heterostructures inevitably contain defects, disorder, and edge roughness, which can influence transport properties. In the present work, our analysis is based on a continuum model in the ballistic regime, where translational invariance along the transverse ($\hat{y}$) direction is assumed. Introducing impurities or rough edges breaks this invariance, making such effects beyond the scope of the current framework. A lattice model would be more appropriate for incorporating disorder and edge roughness in a controlled manner.
VIII Summary and conclusion
@@
Recent theoretical proposals have outlined mechanisms to control the Néel vector in altermagnets [9, 31]. These advances indicate that the phenomena we predict may soon be experimentally testable and could play a role in guiding the development of future spintronic device architectures.